OpenStudy (anonymous):

write an equation in slope-intercept form of the line whose y-intercept is 4and the line is parallel line whose equation is 6+y=5 hat is the slope intercept? i need step by step help please

OpenStudy (anonymous):

i got 2y/2=-6/2x+5/2 so far

OpenStudy (anonymous):

then this equalls y=-3x+4 because i bring down the 4 in the equation as stated in the problem correct?

OpenStudy (anonymous):

So my answer for slope intercept should be y=-3x+4?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

What's your original line?

OpenStudy (anonymous):

6+y=5 is the original along with the 4 as y intercept

OpenStudy (anonymous):

@Brandon77

OpenStudy (anonymous):

okay first get it into y = mx + b form

OpenStudy (kinggeorge):

6+y=5? So y=-1?

OpenStudy (anonymous):

yes, so the slope is 0

OpenStudy (anonymous):

the line you are looking for needs to have a slope of 0 and cross the y axis at 4

OpenStudy (anonymous):

so y = 4

OpenStudy (anonymous):

and so my answer will be y=-3x+4 correct?

OpenStudy (anonymous):

no, your original line had a slope of zero. so your answer should b y=4

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

no prob